# AQORA: A Fast Learned Adaptive Query Optimizer with Stage-Level Feedback for Spark SQL

### v2

@@ VI-B1 Plan Transformation
We observe that applying swap (or lead) operations during a single execution can gradually transform the join plan into a bushy structure. As an illustrative example, consider the swap operation:
$((t_{1}\bowtie t_{2})\bowtie t_{3})\bowtie t_{4}\xrightarrow{\text{swap}((t_{1}\bowtie t_{2}),t_{4})}(t_{1}\bowtie t_{2})\bowtie(t_{3}\bowtie t_{4})$
@@
This example highlights AQORA’s capability to explore bushy join plans through execution-time transformations. By applying the swap operation between the completed $t_{1}\bowtie t_{2}$ subtree and $t_{4}$, the system incrementally constructs a bushy plan during execution.
VI-B2 Join Operator Selection
@@
Upon receiving a broadcast instruction, the planner extension performs a bottom-up traversal of the query plan. It first locates the corresponding join operator within the execution plan hierarchy, then annotates the identified join operator with a BROADCAST hint on the appropriate (left/right) side. This modification instructs Spark SQL to prioritize the broadcast join strategy for the specified join operation.
VII Experiments
In this section, we conduct extensive experiments to validate the effectiveness of AQORA.
VII-A Experiment Setup
VII-A1 Environment
@@
All experiments were conducted on Spark 3.5.4 with 6 executors, each configured with 6 executor cores and 20 GB of executor memory. The decision model was trained and executed for inference on an NVIDIA A40 GPU.
VII-A2 Benchmarks
@@
We selected three standard benchmarks, following prior work [4, 11].
JOB
@@
[35] employs the real-world IMDb dataset with 21 tables related to movies and actors. It includes 113 complex SQL queries generated from 33 templates, each involving joins across 4 to 17 tables.
Extended JOB
@@
[1] (referred to as ExtJOB) is an extension of the JOB benchmark, consisting of 24 queries derived from 12 distinct templates. Compared to the original JOB queries, ExtJOB features entirely different join graphs and predicates, making it a more challenging benchmark[2].
STACK
@@
[5] comprises 10 tables from Stack Exchange datasets. Its 16 query templates generate analytical queries joining 4-12 relations. We excluded templates #4,7 (with arithmetic operations in join conditions) and #9,10 (containing unsupported predicate subqueries)—limitations shared by most existing LQOs. We plan to address these restrictions in the future.
VII-A3 Baselines
We compare AQORA against three state-of-the-art baselines:
Spark SQL
Combined with runtime filters and dynamic join selection, Spark SQL’s default configuration with AQE represents a strong baseline for adaptive query optimization, and it is also the most common industrial practice.
Lero
@@ VII-A3 Baselines
Lero [10] employs a learned comparator to evaluate join plans generated by traditional optimizers under different cardinality estimations, then chooses the best one among them.
AutoSteer
@@
AutoSteer [9] systematically evaluates all available optimization rules in Spark SQL by disabling them to assess their impact on the current plan. It then constructs a collection of rules to disable for performance gains using greedy search.
We adopt the implementations of Lero and AutoSteer from their authors[36, 37]. We exclude other notable optimizers like Bao [5] and Balsa [2] primarily due to their incompatibility with Spark SQL, and we exclude UDAO[27] because it operates in a disjoint optimization space of parameter tuning.
VII-A4 Settings
Data Generation
@@ VII-A4 Settings
For JOB and ExtJOB, we expanded the original IMDb dataset to ten times its original size. For STACK, we utilized the benchmark’s original archived dataset. All datasets were converted to Parquet format. Note that this experimental setup already imposes substantial memory pressure on our testing environment, as an bad query plan can easily lead to out-of-memory errors and execution failure.
Query Generation.
@@
We generated 1,000 training queries using each benchmark’s templates. For each template, randomized predicate conditions were introduced while preserving the original join structure.
For JOB and ExtJOB, we exclusively used their original 113 and 24 queries as their test set, respectively. For STACK, the test set contains 10 additional queries per template generated in the same manner.
Evaluation Approach.
@@
We conduct both static and dynamic evaluations. For the static evaluation, each optimizer is trained on the training set of a given benchmark until convergence, and then its performance is assessed by executing the corresponding test queries. For the dynamic evaluation, we train the optimizer on one dataset and then test it either on the test set of another dataset or on the same dataset after dynamically modifying the underlying data through updates.
Key Metrics
@@
We evaluate performance using three core metrics: (1)End-to-end execution time $C$: total time required to run all test queries; (2)Optimization time $C_{\text{plan}}$: time spent generating optimization decisions; and (3)Raw execution time $C_{\text{execute}}$: time spent on purely executing the query plan. Each query’s execution time is capped at 300 seconds; queries that exceed this limit or fail due to out-of-memory are recorded as 300 seconds.
VII-B Query Performance
Fig. 7 shows the end-to-end execution time, optimization time, and raw execution time for all test queries across different methods.
Figure 7: Query performance on three benchmarks, in seconds. Black numbers outside bars represent end-to-end query execution time, while numbers in dark and light bars represent raw query time and optimization cost respectively.
VII-B1 End-to-End Execution Time
AQORA outperforms all three baselines across the three benchmarks. On the JOB benchmark, AQORA reduces end-to-end execution time by 47.3%, 90.1%, and 60.6% compared with Spark SQL’s default configuration, AutoSteer, and Lero, respectively. For ExtJOB, AQORA achieves reductions of 20.0%, 26.2%, and 38.5% against the same baselines. On STACK, AQORA reduces execution time by 73.3%, 17.2%, and 73.1% compared to Spark SQL’s default, AutoSteer, and Lero, respectively.
These results highlight the advantage of AQORA’s novel idea of optimizing queries during execution, surpassing traditional optimizers and other state-of-art methods.
VII-B2 Optimization Time
Spark SQL’s default optimization strategy incurs no additional overhead, as it directly executes the join order specified in the input SQL text. Consequently, we exclude its optimization time from our comparisons.
For learned methods such as Lero and AutoSteer, the optimization cost primarily stems from generating candidate plans. Lero achieves this by adjusting cardinality estimates at various levels of the query plan, producing over 10 candidate plans per query on average, with each EXPLAIN command taking about 10.1 seconds. AutoSteer, in contrast, generates more than 2 plans per query by enabling or disabling different optimization rules, and each EXPLAIN command takes around 3.3 seconds. This substantial overhead can, in some cases, lead to worse overall performance compared to Spark SQL’s default configuration, even when better query plans are discovered.
AQORA incurs significantly lower optimization costs, as it avoids the need to execute multiple EXPLAIN queries to enumerate candidate plans. Most of its overhead arises from model inference and the execution of Alg. 2, both of which are lightweight. However, AQORA introduces a new type of optimization overhead in distributed engines: modifying partially executed plans at runtime may trigger additional shuffle operations. Despite this, our experiments show the average optimization cost remains below 320 ms per query.
VII-B3 Raw Execution Time
The results show that the raw execution time of AQORA is comparable to that of Lero across all benchmarks. This indicates that modifying query plans during execution can be as effective as generating and evaluating multiple candidate plans from scratch.
The last row of Tab. II summarizes the number and proportion of inferior plans (i.e., out-of-memory or time-limit-exceeded) generated by each method across the three benchmarks. Spark SQL produces a substantial number of inferior plans, particularly on the STACK and JOB benchmarks, primarily due to its limited join optimization capabilities. AutoSteer also produces a number of inferior plans, as its learned optimization strategy tends to favor disabling high-overhead rules. While this can reduce planning time, it often backfires on complex queries. In contrast, AQORA does not generate any inferior plans during test query execution, although a few may arise during training. This highlights the stability and robustness of AQORA’s optimization strategy.
TABLE II: The number of queries showing performance improvement/regression compared with Spark SQL’s default configuration.
JOB ExtJOB STACK
$\delta\in$ Spark Lero AutoSteer AQORA Spark Lero AutoSteer AQORA Spark Lero AutoSteer AQORA
$(0,0.2)$ - 3 1 41 - 4 0 9 - 19 8 22
$(0.2,+\infty)$ - 18 88 20 - 6 0 8 - 51 64 75
(0.2,0)(-0.2,0) - 15 11 41 - 9 8 5 - 19 29 7
$(-\infty,-0.2)$ - 77 13 11 - 5 16 2 - 21 9 6
Failure 10(8.8%) 0 20(17.6%) 0 0 0 0 0 33(30.0%) 0 34(30.1%) 0
VII-B4 Tail Latency
Tail performance is an important indicator of an optimizer’s stability. To provide a comprehensive view, we report the 30th, 60th, 90th, and 99th percentile execution times. Fig. 8 illustrates the end-to-end query latencies of Spark SQL, Lero, AutoSteer, and AQORA at these percentiles after training to convergence. We observe that AQORA consistently outperforms the other methods on all three benchmarks across most percentile levels.
Figure 8: Comparison of execution times at the 30th, 60th, 90th, and 99th percentiles of test queries across three benchmarks for different optimization methods.
On the JOB benchmark, AQORA achieves a 99th percentile latency of 66.86 s, yielding a 99.18% reduction compared to Lero. Notably, the other two baselines fail to produce executable plans at this percentile, underscoring AQORA’s robustness in handling complex queries and its significant advantage in tail latency reduction.
On the Stack benchmark, AQORA consistently outperforms all baselines at the 90th and 99th percentiles. Its raw execution latency at these percentiles is comparable to that of Lero, demonstrating that AQORA can achieve similarly high-quality plans without enumerating multiple candidates.
At lower percentiles, AQORA sometimes underperforms AutoSteer because the latter disables several costly optimization rules. For simpler queries that dominate these percentiles, skipping such rules yields better performance.
VII-B5 Dynamic Evaluation
The first row of Fig. 9 illustrates the performance of AQORA and Lero under data distribution changes. The full IMDb database we use contains data up to 2013. We construct two additional versions, IMDb-1950 and IMDb-1980, by filtering out all movies released after 1950 and 1980, along with their related entries in other tables. IMDb-1950 contains less than 10% of the full IMDb data, while IMDb-1980 retains about 30%[38]. We train the optimizers on these old versions of IMDb and then test them on the complete IMDb database to evaluate their robustness to data evolution. The second row of Fig. 9 shows the results where the optimizers are trained on the training set of JOB and tested on ExtJOB, and vice versa.
Figure 9: The first row shows the performance of AQORA and Lero when trained on IMDb-1950 and IMDb-1980 and tested on the full IMDb dataset. The second row presents the results where the optimizers are trained on the JOB and tested on ExtJOB, and vice versa.
We observe that AQORA consistently outperforms both Spark SQL and Lero under various experimental settings, except for a slight performance drop below Spark SQL when trained on ExtJOB queries using the IMDb-1950 dataset and tested on the full IMDb database. This result demonstrates AQORA’s robustness to dynamic environmental changes. Such robustness mainly stems from two factors: (1) the use of accurate runtime cardinalities enables AQORA to adapt its optimization behavior to the current execution context; and (2) the simplified feature encoding and lightweight network architecture effectively prevent overfitting.
However, both learning-based approaches experience noticeable performance degradation when the execution environment changes. AQORA shows relatively larger drops in raw execution performance. Specifically, Lero’s maximum raw execution degradation across the four settings is 10.80%, 7.86%, 7.53%, and 8.43%, while AQORA’s corresponding drops are 18.04%, 34.93%, 6.99%, and 18.33%. In many cases, AQORA fails to select the optimal plan because the required action was rarely explored during training, whereas Lero can almost always ensure that even an unchosen plan at training at least belongs to its explored set. If AQORA were trained using an offline RL algorithm equipped with a comprehensive experience buffer, this limitation might be mitigated.
VII-C Case Study
In this section, we analyze the results on a per-query basis to find out what AQORA actually achieves in optimizing query plans. In particular, we compare the end-to-end query execution time of learning-based methods with those generated by Spark SQL’s native optimizer. For a query plan $p$ executed by a learning-based method, we calculate its improvement or regression as $\delta=(C_{\text{Spark}}(p)-C_{\text{LQO}}(p))/C_{\text{Spark}}(p)$. Tab. II summarizes the distribution of $\delta$ for each method.
VII-C1 Query Improvement Analysis
As shown in Fig. 10, AQORA achieves up to 89.81%, 69.33%, and 91.83% performance improvement on individual queries from the JOB, ExtJOB, and STACK benchmarks, respectively.
@@ VII-C1 Query Improvement Analysis
Figure 10: The top 10 queries improved by AQORA in each benchmark compared with Spark SQL’s default configuration.
@@
For example, in JOB query Q16a, AQORA not only enabled the CBO but also refined the plan by prioritizing the join between table title and table movie_keyword, significantly reducing the execution time from over 300 s to just 30.56 s, demonstrating AQORA’s ability to refine CBO-generated plans by improving join orders while balancing optimization overhead and performance gains.
@@
VII-C2 Query Regression Analysis
@@
While AQORA delivers substantial improvements on many queries, we also observe performance regressions on a small subset. Minor regressions are due to the optimization overhead introduced by AQORA.
@@
Significant regressions stem from limitations of AQORA’s current learning framework, which occasionally leads to suboptimal plan choices (e.g., JOB Q28b). To mitigate this, we introduced several restrictions during training, such as limiting the number of candidate lead tables and capping the maximum number of action steps. These constraints help guide the agent toward more stable decisions and reduce exploration noise. However, such regressions are not entirely avoidable and remain an area for future improvement.
VII-C3 Proportion of bushy plans
@@
Here we analyze the proportion of test queries whose final executed plans adopt a bushy join structure. In our experiments, 22 (19.5%) queries in JOB, 5 (20.1%) in ExtJOB, and 30 (27.3%) in STACK ultimately execute with bushy plans, enabled by the application of CBO actions and leading actions as described in Sec. VI-B1. These results demonstrate that AQORA is capable of effectively exploring bushy join plans to improve query performance.
VII-D Ablation Study

### v3

@@ 5.2.1. Plan Transformation
We observe that applying swap or lead action during query execution can gradually transform the join plan into a bushy structure. As an illustrative example, consider the swap action in Figure  5:
Figure 5. A swap between $(t_{1}\bowtie t_{2})$ and $t_{4}$ can turn the left-deep join plan into a bushy join plan.
@@
By applying the swap action between the completed $t_{1}\bowtie t_{2}$ subtree and $t_{4}$, the planner extension incrementally constructs a bushy plan during execution.
5.2.2. Join Operator Selection
@@
Upon receiving a broadcast action, the planner extension performs a bottom-up traversal of the query plan. It first locates the corresponding join operator within the execution plan hierarchy, then annotates the identified join operator with a BROADCAST hint on the appropriate (left/right) side. This modification instructs Spark SQL to prioritize the broadcast join strategy for the specified join operation.
5.2.3. After Plan Injection
After the optimization actions are applied, the modified plan is handed back to the AQE execution engine for continued execution. LQRS does not restart query execution or invalidate completed query stages. Instead, the injected modifications only affect the remaining unexecuted portion of the plan.
In addition, the planner extension benefits from AQE’s built-in plan validation mechanisms. While LQRS does not rely on Spark’s internal plan evaluation or cost comparison and always adopts the plan produced by its optimization actions, AQE still enforces a set of critical safety checks before execution is resumed. In particular, the AQE engine guarantees the semantic equivalence of the modified plan with respect to the original one and ensures that the injected plan can be correctly compiled and executed by the underlying engine. As a result, the planner extension does not need to handle low-level concerns such as semantic equivalence checking, plan code generation, or execution resumption.
5.3. Compatibility with Other Systems
While the planner extension in LQRS is implemented on top of the Spark AQE framework, its design principles are applicable to other database systems. A key challenge in generalizing such extensions lies in identifying points at which execution can be paused or re-optimized with acceptable overhead. Spark SQL naturally provides such boundaries through query stages, which explicitly materialize intermediate results and support execution suspension and re-optimization. For database systems such as Flink SQL in batch mode, Presto, and Snowflake, which have similar execution models, natural synchronization points arise where intermediate results are produced. In these systems, planner extensions can likewise use these materialization boundaries to perform runtime optimization.
For other systems that do not provide such exploitable execution boundaries, we need to identify points where execution can be safely paused or re-optimized. In PostgreSQL, We observe that pipeline-breaking operators, such as hash and sort operators, can be synchronization points because they inherently materialize intermediate results and interrupt tuple pipelines. DuckDB employs a vectorized execution engine, yet certain operators still require blocking behavior, creating implicit plan boundaries where runtime statistics can be observed. These observations indicate that even without explicit AQE support, our planner extension can use existing materialization or pipeline breakers for re-optimization.
6. Experiments
6.1. Experiment Setup
6.1.1. Environment
@@
All experiments were conducted on Spark 3.5.4 with 6 executors, each configured with 6 cores and 20 GB memory. The decision model was trained and executed for inference on an NVIDIA A40 GPU.
6.1.2. Benchmarks
@@
We selected four standard benchmarks, following prior work  (Chen et al., 2024a; Yu et al., 2022; Chen et al., 2024b):
@@
JOB (Leis et al., 2015): JOB employs the real-world IMDb dataset with 21 tables related to movies and actors. It includes 113 complex SQL queries generated from 33 templates, each of them involving joins across 4 to 17 tables.
@@
Extended JOB (referred to as ExtJOB(Marcus et al., 2019): ExtJOB is an extension of the JOB benchmark, consisting of 24 queries derived from 12 distinct templates. Compared to the original JOB queries, ExtJOB features entirely different join graphs and predicates, making it a more challenging benchmark (Yang et al., 2022).
@@
STACK (Marcus et al., 2021): STACK comprises 10 tables from Stack Exchange datasets. Its 16 query templates generate analytical queries joining 4-12 relations. We excluded templates #4,7 (with arithmetic operations in join conditions) and #9,10 (containing unsupported predicate subqueries)—limitations shared by most existing LQOs. We plan to address these restrictions in the future.
TPC-H (1): TPC-H is designed to test the response time of database systems to complex queries. It includes eight tables with the scale factor set to 100 in our experiments. In our evaluation, we only consider the query templates #2, 3, 5, 7, 8, 9, 10 following  (Chen et al., 2024b). We exclude other templates, as they are too simple (only have one or two tables) or contain predicate subqueries.
6.1.3. Baselines
We compare LQRS against four baselines:
Spark SQL: Spark SQL with its default AQE mechanism.
@@ 6.1.3. Baselines
Lero (Chen et al., 2024b): Lero employs a learned comparator to evaluate join plans generated by Spark SQL’s optimizer under different cardinality estimations, then chooses the best one among them.
@@
AutoSteer (Anneser et al., 2023): AutoSteer systematically evaluates all available optimization rules in Spark SQL by disabling them to assess their impact on the current plan. It then constructs a collection of rules to disable for performance gains using greedy search.
SSA: QuerySplit (Zhao et al., 2023) is a query re-optimization algorithm that proactively decomposes a query into multiple subqueries at the logical plan level and iteratively selects subqueries to execute based on their estimated cost and output cardinality. In Spark SQL, query plans is naturally partitioned into query stages at shuffle boundaries. Therefore, we only re-implement the subquery selection component of QuerySplit, namely SSA, within Spark SQL and use it as a query re-optimization baseline for comparison.
6.1.4. Settings
For static evaluation, we refer to the experimental settings of Bao (Marcus et al., 2021) and Lero (Chen et al., 2024b), while for dynamic evaluation, we refer to the settings used in Robust-MSCN (Negi et al., 2023):
@@ 6.1.4. Settings
Data Generation. For JOB and ExtJOB, we expanded the original IMDb dataset to ten times its original size. For STACK, we utilized the benchmark’s original archived dataset. For TPC-H, we generated the dataset using the official data generation tool with a scale factor of 100. All datasets were converted to Parquet format. Note that this experimental setup already imposes substantial memory pressure on our testing environment, as a bad query plan can easily lead to out-of-memory errors and execution failure.
@@
Query Generation. We generated 1,000 training queries using each benchmark’s templates. For each template, randomized predicate conditions were introduced while preserving the join structure. For JOB and ExtJOB, we exclusively used their original 113 and 24 queries as their test set, respectively. For STACK and TPC-H, the test set contains 10 additional queries per template generated in the same manner.
@@
Evaluation Approach. We conduct both static and dynamic evaluations. For the static evaluation, each optimizer is trained on the training set of a given benchmark until convergence, and then its performance is assessed by executing the corresponding test queries. For the dynamic evaluation, we train the optimizer on one dataset and then test it either on the test set of another dataset or on the same dataset after dynamically modifying the underlying data through updates.
@@
Key Metrics. We evaluate performance using End-to-end execution time: total time required to run all test queries. Each query’s end-to-end execution time is capped at 300s; queries that exceed this limit or fail due to OOM error are recorded as 300s.
6.2. Query Performance
Table 3 reports the end-to-end execution time across all benchmarks.
Table 3. Overall performance on all benchmarks.
Time (s) JOB ExtJOB STACK TPC-H
Spark SQL 5821.4 694.0 13149.1 2954.0
Lero 33211.9 752.9 4231.6 2625.0
AutoSteer 7793.3 903.2 13026.1 3010.2
SSA 3124.1 651.0 7934.2 4895.7
LQRS 3069.5 555.5 3504.3 2471.3
LQRS consistently achieves the best performance on all four workloads, outperforming all baseline methods. Compared with Spark SQL, LQRS reduces the total execution time by 16.3%–73.3% across four benchmarks. These results highlight the advantage of LQRS’s novel idea of allowing learned optimization decisions to be deferred to execution time and guided by actual runtime observations.
Figure 6 illustrates the performance improvement and regression of each query on four tested benchmarks with Lero, AutoSteer, SSA, and LQRS w.r.t the corresponding end-to-end query execution latency in Spark SQL. We omit cases where Spark SQL chooses a similar plan (with a runtime difference << 5s) to LQRS and other baseline methods. The key observations are as follows:
Figure 6. Latency improvement and regression of each query in Lero, AutoSteer, SSA, and LQRS w.r.t Spark SQL
  • Compared with Lero, our performance improvements exhibit similar trends, but we observe fewer regressions, especially on JOB. This indicates that LQRS can leverage runtime observations to obtain execution plans of comparable quality to Lero, while avoiding the extensive plan enumeration and evaluation overhead that Lero incurs, highlighting the effectiveness of LQRS.
  • Compared with AutoSteer, our improvements are substantially larger. AutoSteer’s learned optimization strategy tends to favor disabling high-overhead rules, which often backfires on complex queries. In contrast, LQRS does not generate any inferior plans during test query execution, highlighting the robustness of its optimization strategy.
  • Compared with SSA, our performance gains are more substantial and stable, with fewer large-scale regressions observed on TPC-H. This is because our reinforcement learning–based decision process is more comprehensive: it evaluates feedback at each step of the actor and considers the long-term value of partially executed plans. As a result, LQRS does not greedily lead low-cost leaf nodes forward, but instead balances immediate costs with long-term benefits, highlighting the intelligence of its decision-making process.
6.3. Case Study
In this section, we analyze the results on a per-query basis to find out what LQRS actually achieves in optimizing query plans.
6.3.1. Query Improvement Analysis
@@
Figure  7 shows the top 10 queries improved by LQRS in each benchmark compared with Spark SQL. LQRS achieves up to 89.81%, 69.33%, 91.83%, and 55.21% performance improvement on the JOB, ExtJOB, STACK, and TPC-H benchmarks, respectively.
Figure 7. The top 10 queries improved by LQRS in each benchmark compared with Spark SQL.
@@
For example, in JOB query Q16a, LQRS not only chooses the cost-based initialization strategy but also refines the plan by prioritizing the join between table t and table mk, significantly reducing the execution time from over 300s to just 30.56s, demonstrating LQRS’s ability to refine CBO-generated plans by improving join orders while balancing optimization overhead and performance gains.
@@
6.3.2. Query Regression Analysis
@@
While LQRS delivers substantial improvements on many queries, we also observe performance regressions on a small subset. Minor regressions are mainly due to the optimization overhead introduced by LQRS.
Significant regressions in query performance mostly stem from inherent limitations of LQRS’s current learning framework, which occasionally results in suboptimal plan selections. For instance, JOB Q8d experiences performance degradation because the table t was incorrectly selected as the lead table, causing a substantial increase in intermediate data movement. As a result, the shuffle size during query execution nearly tripled compared to the baseline plan, significantly impacting runtime.
@@
To mitigate such regressions, we introduced several training-time constraints, including limiting the number of candidate lead tables and capping the maximum number of action steps. These restrictions guide the agent toward more stable and predictable decisions while reducing exploration noise. Nevertheless, regressions of this kind cannot be entirely avoided under the current framework and represent an important avenue for future improvement.
6.3.3. Proportion of bushy plans
@@
Here, we analyze the proportion of test queries whose final executed plans adopt a bushy join structure. In our experiments, 22 (19.5%) queries in JOB, 5 (20.1%) in ExtJOB, and 30 (27.3%) in STACK ultimately execute with bushy plans, enabled by the plan transformation actions as described in Section 5.2.1. These results show that LQRS is capable of effectively exploring bushy join plans to improve query performance.
6.4. Dynamic Evaluation
Figure 8 evaluates the robustness of LQRS and Lero under data distribution and workload changes. We chose Lero due to its relatively strong adaptability among all the baselines.
For data distribution changes (Figure 8(a–b)), we construct two historical variants of IMDb—IMDb-1950 and IMDb-1980—by retaining only movies released no later than 1950 and 1980, respectively, together with their related records. These variants contain less than 10% and about 30% of the full IMDb dataset (Negi et al., 2023). The optimizers are trained on these historical datasets and evaluated on the full IMDb to assess robustness to distribution shifts. For workload changes (Figure 8(c–d)), the optimizers are trained on either JOB or ExtJOB and evaluated on both workloads, enabling cross-workload generalization analysis.
Figure 8. Performance of LQRS and Lero under data distribution and workload changes. (a–b) Optimizers trained on older IMDb snapshots (1950, 1980) and tested on the full database. (c–d) Optimizers trained on either JOB or ExtJOB workloads and tested on both JOB and ExtJOB.
We observe that LQRS consistently outperforms both Spark SQL and Lero under various experimental settings, except for a slight performance drop below Spark SQL when trained on ExtJOB queries using the IMDb-1950 dataset and tested on the full IMDb database. This result demonstrates LQRS’s robustness to environmental changes. Such robustness mainly stems from two factors: (1) the use of runtime observation enables LQRS to adapt its optimization behavior to the current execution context; and (2) the simplified feature encoding architecture effectively prevents overfitting.
However, both learning-based approaches suffer performance degradation under execution environment changes, with LQRS exhibiting relatively larger drops in raw execution performance. Across the four settings, Lero’s maximum degradation ranges from 7.53% to 10.80%, whereas LQRS experiences more severe drops of up to 34.93%. In many cases, LQRS fails to select the optimal plan because the required action was rarely explored during training, whereas Lero can almost always ensure that even an unchosen plan at training at least belongs to its explored set. If LQRS were trained using an offline RL algorithm equipped with a comprehensive experience buffer, this limitation might be mitigated.
6.5. Ablation Study